### v1

# Population demographics of white dwarf binaries with intermediate separations: Gaia constraints on post-AGB mass transfer

Abstract
Astrometry from the Gaia mission has revealed a large population of white dwarf (WD) + main sequence (MS) binaries with periods of $100-1000\,$d. Such systems have separations intermediate to predictions from standard binary evolution scenarios, challenging models of binary interaction and mass transfer. Because the selection function of Gaia astrometric catalogs is complex, the underlying population demographics of WD+MS binaries remain imperfectly understood. We present a forward-model of the au-scale WD+MS binary population probed by Gaia that begins with a realistic binary population and incorporates a full model of Gaia mock observations and astrometric model fitting, as well as cuts employed in producing the Gaia astrometric catalog and selecting WD+MS binary candidates. Our model allows us to constrain the intrinsic population demographics of intermediate-separation WD+MS binaries. The inferred period distribution is close to flat, with ${\rm d}N/{\rm d}P_{{\rm orb}}\propto P_{{\rm orb}}^{-0.25}$, while the WD mass distribution is sharply peaked at $0.6\,M_{\odot}$. We model the formation of au-scale WD+MS binaries as the result of interaction when the WD progenitor is an AGB star. Explaining the mass distributions of both components requires two key assumptions: (1) the mass growth of the WD is terminated when its AGB progenitor overflows its Roche lobe, and (2) post-AGB binaries binaries remain wide only if the accretor-to-donor mass ratio exceeds a critical threshold, $q_{\rm crit}\approx 0.4$. Systems with more unequal mass ratios likely shrink to shorter periods and become classical post-common envelope binaries or merge. The model implies that $\sim 1\%$ of solar-type stars have WD companions with periods of $100-1000\,$d, consistent with complementary constraints from self-lensing binaries.
Binary stars (154) — White dwarf stars (1799) — Astrometry (80)
facilities: Gaiasoftware: gaiamock (El-Badry et al., 2024a), stam (Hallakoun & Maoz, 2021)
1 Introduction
Close white dwarf (WD) + main sequence (MS) binaries are end products of binary interactions that occurred when the WD progenitor was a giant. They are therefore useful probes of the physics of mass transfer (MT) processes which play an important role in the evolution of many types of binaries.
The third data release (DR3) of the Gaia mission introduced the non-single stars (NSS) catalog, which included orbital solutions for an unprecedented number of binaries. In particular, it contains astrometric orbital solutions for over 160,000 sources, leading to the discovery of new populations of luminous stars orbited by black holes, neutron stars, and WDs (El-Badry et al., 2023a, b; Chakrabarti et al., 2023; El-Badry et al., 2024b; Shahaf et al., 2023, 2024). To identify astrometric binaries likely to contain compact objects, Shahaf et al. (2019) developed a “triage” scheme. It rests on a quantity called the astrometric mass ratio function (AMRF), which depends only on observables from Gaia and is larger for systems hosting dark companions than for systems hosting luminous companions. Using this technique, Shahaf et al. (2024) identified a sample of over 3000 high-probability WD+MS binaries with orbital periods ranging from $\sim 100-1000\,$days, corresponding to AU-scale separations. While post-interaction binaries with comparable separations were known before Gaia (e.g. Kawahara et al., 2018; Masuda et al., 2019; Oomen et al., 2018; Jorissen et al., 2019; Escorza et al., 2019), the scale of the population was not appreciated because most photometric and spectroscopic searches are more sensitive to short than to long orbital periods.
The abundance of these AU-scale WD+MS binaries was unexpected, because standard binary evolution models predicted few such systems to exist. While the binaries’ separations are too close to have avoided interaction when the WD progenitor was on the asymptotic giant branch (AGB) – implying they are post-MT systems – their orbits are significantly longer than predicted by models of dynamically unstable MT (i.e. common envelope evolution, CEE; e.g. Paczynski 1976; Ivanova et al. 2013). An alternative possibility is that the systems are products of stable mass transfer, but current evolutionary models predict unstable mass transfer for the systems’ inferred progenitor mass ratios (e.g. Yamaguchi et al., 2024b). Therefore, this sample highlights the imperfect current understanding of MT processes and motivates detailed studies.
While several works have attempted to model the evolution of wide post-interaction WD+MS binaries (e.g. Belloni et al., 2024; Yamaguchi et al., 2024b), a population-level study requires an understanding of the selection function. This is made challenging by the fact that Gaia DR3 does not contain epoch astrometry or orbital solutions for all sources. Instead, each source goes through a multi-step pipeline, where cuts are made to remove potentially spurious solutions and where a full astrometric binary model is only attempted if simpler models do not provide a sufficiently good fit to the data. This means that while the published orbits are largely reliable (e.g. Yamaguchi et al., 2024a), many true binaries do not make it into the NSS catalog. Inference of population-level properties of the WD+MS binary sample, such as period, eccentricity, and mass ratio distributions, has thus far remained limited by the poorly-understood selection function of the NSS catalog.
Recently, El-Badry et al. (2024a) (E24 henceforth) used a forward-model of Gaia epoch astrometry and astrometric solution processing to approximate the Gaia astrometric orbit selection function. Starting with a model of the Galactic binary star population, they produced synthetic Gaia epoch astrometry for each source. They then ran the synthetic astrometry through the same cascade of astrometric models used in producing the NSS catalog. They showed that the distributions of orbital and stellar parameters of the binaries in the resulting mock catalog are in good agreement with those of the real catalog, implying that their selection function is a good approximation of reality.
In this work, we extend the work of E24, focusing on the WD+MS binary population. We self-consistently forward-model the creation of the Gaia DR3 astrometric orbit sample, the selection of WD+MS binary candidates via the AMRF, and potential contamination of the sample from hierarchical triples, tuning the assumed underlying WD+MS binary population to match the final observed sample from Shahaf et al. (2024). In doing so, we constrain the occurrence rate and population demographics of WD+MS binaries.
The rest of this paper is organized as follows: In Section 2, we summarize the steps taken by E24 to create a mock Gaia NSS catalog and describe our addition of hierarchical triples to their model. In Section 3, we provide an overview of the triage technique employed by Shahaf et al. (2024) to construct the observed sample of WD+MS binaries using orbital solutions from Gaia DR3. We also describe its application to our simulated binaries. In Section 4, we compare properties of the resulting mock sample of WD+MS binaries obtained using the original model from E24 to those of the observed sample. Based on this, in Section 5, we describe an improved model that better matches the data. In Section 6, we compare the model’s predictions to data and discuss completeness fractions and the underlying parameter distributions of WD+MS binaries. Finally, we conclude in Section 7.
2 Forward model of the Gaia DR3 NSS catalog
We begin by summarizing the E24 model of Gaia astrometric orbit catalogs. We refer readers to E24 for a more detailed description.
2.1 Initial binary population
[…]
Since Galaxia does not include binaries, our model uses COSMIC (Breivik et al., 2020) to generate a zero-age binary population with parameter distributions from Moe & Di Stefano (2017). Binaries are matched to stars in the Galaxia population by their primary masses and placed at the same position. Binary orientations and phases are assigned randomly. G-band magnitudes are calculated from MIST isochrones (Morton, 2015; Choi et al., 2016) for each binary component. We calculate extinctions using the combined19 dust map in the mwdust package (Bovy et al., 2016), which combines dust maps from Drimmel et al. (2003); Marshall et al. (2006); Green et al. (2019).
2.2 Formation of compact objects
Binaries with primaries with initial masses above $8\,M_{\odot}$, which have terminated their evolution as black holes or neutron stars, are removed under the assumption that most of them are broken up or shrink to very tight orbits via CEE. Stars with initial masses below $8\,M_{\odot}$ are assumed to end their lives as WDs. E24 predicted WD masses using the empirical initial-final-mass relation (IFMR) of Weidemann (2000), which is calibrated to observations of single WDs. In this work, we explore a different IFMR that accounts for binary interaction (Section 5.2).
2.3 Post-interaction separations
E24 assumed that binaries with initial separations $<2\,$AU experience mass transfer which efficiently shrinks and circularizes their orbits such that they end up in 1-day orbits with zero eccentricities which are not detectable with astrometry. Meanwhile, they assumed that 10% of orbits with initial separations between $2$ and 666\,6AU end up in wider orbits with final separations half of their initial separations and moderate final eccentricities randomly selected between 0 and 0.2 (the remaining 90% end up in 1-day circular orbits). The 222\,2AU boundary was chosen to roughly distinguish between systems that interact with donors on the red giant branch (RGB) versus the AGB. This was motivated by recent studies of the Gaia WD+MS binary sample which have found that it is possible to from relatively wide post-common envelope binaries (PCEBs) if mass transfer begins during a thermal pulse of an AGB donor (e.g. Belloni et al., 2024; Yamaguchi et al., 2024a). These assumptions regarding the orbital evolution due to MT are modified in this work (Section 5.1).
E24 did not model the formation of low-mass helium WDs that are formed through stable mass transfer from donors on the first giant branch (e.g. Rappaport et al., 1995; Marsh et al., 1995; Brown et al., 2016; Li et al., 2019; El-Badry & Rix, 2022; Garbutt et al., 2024), and we also do not attempt to track such binaries in this work. Although low-mass WDs formed via stable mass transfer are common in samples of WD binaries selected via UV excess, they are almost entirely absent in the AMRF sample, which is only sensitive to systems with $M_{\rm WD}/M_{\star}\gtrsim 0.6$ (El-Badry, 2024).
2.4 Mock observations
We obtain mock observations for each source based on the Gaia scanning law using the Gaia Observation Forecast Tool (GOST) [^1] https://gaia.esac.esa.int/gost, and model the effect of binarity on the epoch astrometry following Lindegren (2022). We adopt an empirical noise model based on the residuals of DR3 epoch astrometry for well-behaved sources from Holl et al. (2023). Finally, we process the resulting 1D astrometric measurements through a pipeline similar to that described in Halbwachs et al. (2023) to produce a mock NSS catalog. In Figure 1, we show several simulated WD+MS binaries, along with their mock observations. The leftmost panel shows a typical WD+MS binary in the AMRF sample with $P_{\rm orb}=600\,$d at a distance of 300300300\,300pc. This enters the mock NSS catalog with a full astrometric orbital solution. In the middle panel is a case of an orbit that receives an orbital solution but is excluded from the final catalog as it does not satisfy the quality cut imposed on the relative parallax uncertainty to remove potentially spurious solutions ($\varpi/\sigma_{\varpi}<20000/P_{\rm orb}$; Halbwachs et al. 2023). Finally, the orbit on the right panel, placed at a larger distance, only receives a 7-parameter acceleration solution and therefore also does not enter the NSS catalog.
The results in E24 can be reproduced using their publicly available code [^2] https://github.com/kareemelbadry/gaiamock.
Figure 1: Examples of orbits and their mock observations. In all cases, we consider a $1.0\,M_{\odot}$ luminous primary and a dark $0.6\,M_{\odot}$ secondary in a slightly eccentric ($e=0.1$) orbit. Different orbital periods and distances are considered in each panel. The ratio of semi-major axis and parallax to their respective errors are quoted on the bottom of each panel; quality cuts on these parameters determine whether each binary receives an orbital solution. The leftmost panel is most representative of a typical system in the AMRF sample and receives an orbital solution. The middle one is excluded from the NSS catalog due to having $\varpi/\sigma_{\varpi}$ that is too low, despite the orbit being fairly well constrained. The rightmost orbit gets an acceleration solution and is therefore not fit with an orbital solution.
2.5 Triples
Because an inner binary of two MS stars has a higher mass-to-light ratio than a single star, hierarchical triples are the most important possible contaminant for an AMRF-selected WD sample (Section 3). To study their contribution to the final sample, we modify the E24 model to include triples. We assign an outer tertiary to a subset of the zero age binary population with probability that depends on the primary mass. The triple fraction as a function of primary mass is obtained from Table 1 of Offner et al. (2023). Since we are assigning triples to existing binaries, we use the fraction of triples and higher-order multiples in all multiple star systems (in the terminology adopted by Offner et al. 2023, this is THF/MF). For simplicity, in this step, we define the primary as the most massive component of the inner binary. While this is not exact (as the outer tertiary may end up most massive), Figure 2 shows that this procedure results in a triple fraction close to the observational constraints of Offner et al. (2023).
Figure 2: Fraction of the simulated zero-age binary population that are in hierarchical triples as a function of the true primary mass (i.e. maximum mass of the two/three components in the system). In blue is the inputted triple fraction from Offner et al. (2023).
For systems selected to be in triples, we draw masses of the outer tertiary assuming a uniform distribution of the mass ratio of the outer tertiary over the inner binary. The eccentricities of the outer orbits are sampled from a thermal distribution, and the periods from a log-normal distribution with a mean in log($P/\mathrm{d}$) of 4.54 and standard deviation of 2.4 (Tokovinin, 2014). This process is repeated until each system meets the stability criteria for hierarchical triples from Naoz et al. (2013) and Mardling & Aarseth (2001) (as summarized in Section 2.3 of Shariat et al. 2023).
For the astrometric processing, we assume that if the angular size of the outer orbit (= semi-major axis/distance) is more than 1” on the sky, the outer tertiary is resolved and it will receive its own 5-parameter solution (Nagarajan & El-Badry, 2024). Therefore, the inner binary of such a system is processed independently. If the size of the outer orbit is less than 1” and its orbital period is less than 1000 d, the system is modeled as a binary made up of the inner binary and outer tertiary. The total mass and luminosity of the inner binary is taken as the mass and luminosity of one component. Such a system may receive an astrometric binary solution which can contaminate the AMRF sample. We discard triples with outer orbital periods greater than 1000 d with luminous tertiaries as these periods are too long for the outer orbit to be constrained in Gaia DR3, while the presence of the tertiary likely disturbs the astrometric measurements of the inner orbit (e.g. Lindegren, 2022). We also neglect systems where both components of the inner binaries have evolved into WDs, since our modeling is unlikely to reliably track the mass transfer that preceeds the second WD’s formation.
As we discuss in Section 6, the color excess cut ultimately removes the vast majority of triples, meaning that the final WD+MS binary sample is relatively insensitive to our assumptions about the triple population.
For most of the subsequent analysis, we only consider triples and WD+MS binaries, excluding all other binaries which host two luminous companions as these are almost entirely removed by the cut on the AMRF and thus do not enter the final simulated WD+MS binary sample.
In Figure 3, we show the distributions of astrometric mass function, $f_{\rm m,ast}$, of the orbital solutions in the mock NSS catalog resulting from our fiducial model (described in Section 5) as well as the true NSS catalog published in Gaia DR3. To isolate MS primaries and exclude giants, we only plot sources in the true catalog which are also present in the binary_masses catalog and consider simulated binaries with primary radii $<2\,R_{\odot}$. $f_{\rm m,ast}$ is given by
$f_{\rm m,ast}=\left(\frac{\alpha}{\varpi}\right)^{3}\left(\frac{P_{\rm orb}}{1\mathrm{yr}}\right)^{-2}.$ (1)
It is useful for identifying both spurious astrometric solutions (Halbwachs et al., 2023) and systems with compact object companions. Here, we also include non- WD+MS binaries from the E24 whose properties are minimally affected by the changes made in this work. Looking at the simulated binaries, most of the low-$f_{\rm m,ast}$ solutions are non-WD+MS binaries while the high-$f_{\rm m,ast}$ solutions are triples. WD+MS binaries lie in between, with a median value of $0.07$. In total (not excluding giants), there are 187,000 orbital solutions in our mock catalog which is $12\%$ more than the 168,000 in the true catalog. For comparison, the original E24 catalog contained $18\%$ fewer sources than the true catalog. The difference is primarily attributed to the increased number of WD+MS binaries in our model. As we will show (Section 6), the typical WD mass in the sample is lower than assumed by E24. As a result, a smaller fraction of WD+MS binaries enter the AMRF sample than assumed in their work, and so a larger underlying WD+MS binary population is required to explain the same AMRF sample. We see that the distributions are in good agreement to one another, though it is more peaked for the mock catalog. WD + MS binaries make up $28\%$ of sources in our mock catalog, and $\sim 7\%$ of these make it into the final sample.
Figure 3: Distributions of the astrometric function, $f_{\rm m,ast}$, for the orbital solutions in the simulated fiducial (green) and true (blue) NSS catalogs. Contributions from the different components in the simulated catalog are shown with unfilled histograms. We see that the total number of orbital solutions as well as the $f_{\rm m,ast}$ distribution of the mock and true catalog are in rough agreement with each other.
3 AMRF sample of WD+MS binaries
We apply the same cuts to mock binary catalogs as were applied to the NSS catalog by Shahaf et al. (2024). We first calculate the AMRF, $\mathcal{A}$:
$\mathcal{A}=\frac{\alpha}{\varpi}\left(\frac{M_{1}}{M_{\odot}}\right)^{-1/3}\left(\frac{P_{\rm orb}}{\mathrm{d}}\right)^{-2/3}$ (2)
where $\alpha$ is the angular photocentric semi-major axis, $\varpi$ is the parallax, $M_{1}$ is the primary mass, and $P_{\rm orb}$ is the orbital period. This can be rewritten in terms of the mass ratio, $q$, and G-band flux ratio, $\mathcal{S}$, of the system:
$\mathcal{A}=\frac{q}{(1+q)^{2/3}}\left(1-\frac{\mathcal{S}(1+q)}{q(1+\mathcal{S})}\right)$ (3)
where the ratios are defined as the photometric secondary over primary.
Gaia provides constraints on all parameters in equation 2. Shahaf et al. (2024) obtained primary mass estimates from the binary_masses catalog (Gaia Collaboration et al., 2023), while they calculated $\alpha$ from the Thiele-Innes coefficients reported in the nss_two_body_orbit catalog.
For our simulated systems, we use the true primary masses from our model. This implicitly assumes that the masses reported in the binary_masses catalog are accurate and do not suffer systematic biases. We assume a fixed uncertainty of $0.06\,M_{\odot}$, which is the average value of the reported errors for the sources in the observed sample. We remove simulated binaries with primary radii $>2\,R_{\odot}$, as evolved stars are not included in the binary_masses catalog and are thus also excluded by Shahaf et al. (2024).
Using theoretical mass-luminosity relations for a range of primary masses, Shahaf et al. (2024) calculate boundaries in the $\mathcal{A}-M_{1}$ space to define several classes corresponding to the most likely types of secondaries: (1) a single MS star, (2) an inner binary of two MS stars (i.e. hierarchical triple), and (3) a dark companion. As a starting point, they isolated systems in the latter two classes (i.e. inconsistent with hosting a single MS companion), which they referred to as the non-class I sample.
For further details regarding the application of this technique to the orbits from the real NSS catalog, we refer readers to Shahaf et al. (2023, 2024).
3.1 Color excess cut
Figure 4: The four panels show 2D binned plots of the color excess against various stellar and orbital parameters of all sources in the non-class I catalog of Shahaf et al. (2024).
To reduce contamination from hierarchical triples, Shahaf et al. (2024) removed systems from the non-class I sample that are redder than expected for a single luminous companion. They did this by calculating the “color excess” (CE) of each source:
$\Delta(B-I)=(B-I)_{\rm observed}-(B-I)_{\rm expected}$ (4)
where $(B-I)_{\rm observed}$ and $(B-I)_{\rm expected}$ are the de-reddened observed and theoretically expected $B-I$ colors.
Shahaf et al. (2024) took observed B, V, and I band magnitudes generated using the BP/RP spectra from the Gaia synthetic_photometry_gspc catalog and used 3D dust extinction maps (Green et al., 2019; Lallement et al., 2019) to correct for reddening. To calculate $(B-I)_{\rm expected}$, they interpolated PARSEC isochrones with a fixed age of 2 Gyrs on metallicity and the absolute V band magnitude using the stam333github.com/naamach/stam package (Hallakoun & Maoz, 2021). They obtained metallicity estimates from Zhang et al. (2023) who created an empirical forward model of the Gaia XP spectra to obtain stellar atmospheric parameters for 220 million sources. For the $\sim 20\%$ of sources flagged to have unreliable stellar parameters by Zhang et al. (2023), they set [Fe/H] = 0.0 with an uncertainty of 0.25 dex. For an in-depth description of this process, consult Section 2.2 of Shahaf et al. (2024).
For each source, Shahaf et al. (2024) calculated $10^{4}$ realizations of the color excess, each time drawing observables from uncorrelated Gaussian distributions with standard deviations set to their errors. Sources where the fraction of realizations with $\Delta(B-I)>0$ exceeded 64% (corresponding to the value above which the infrared excess is greater than than 1%) were removed [^4] An error in the code used by Shahaf et al. (2024) led this threshold to be 56% in the first-published catalog. This has since been corrected and we compare our models to this updated catalog (Shahaf et al., 2025)., leaving behind $\sim 3000$ “high probability” WD+MS binaries in the final “no-color excess” (NCE) sample.
For the simulated population, we predict the B, V, and I magnitudes of each system by interpolating on PARSEC isochrones and assuming the sources’ true age and metallicity. Our grid of isochrones have log(age) ranging from 6.5 to 10.5 in steps of 0.1 dex and [M/H] from -2.0 to 1.0 in steps of 0.1 dex. We calculate the observed apparent magnitudes from the total flux of all components which are then extincted/reddened according to the true distance to the source. We use the dustmaps (Green, 2018; Green et al., 2019) and mwdust python package (Bovy et al. 2016; Drimmel et al. 2003; Marshall et al. 2006; Green et al. 2019) to predict extinctions. We calculate the uncertainties in these apparent magnitudes via interpolation of an exponential fit to the flux errors against the V band fluxes of sources in the observed non-class I sample. We also fit residuals to the best-fit curve with an exponential, which we used to add noise to the interpolated errors, drawn from a normal distribution with a standard deviation given by the residual model. Finally, we convert the apparent magnitudes to “observed” absolute magnitudes using the measured parallaxes from the mock astrometric fits.
We calculate $(B-I)_{\rm expected}$ using the same code as Shahaf et al. (2024), similarly assuming a fixed age of 2 Gyr. For the majority of sources, we use the true values of [Fe/H] generated by Galaxia and assume a fixed uncertainty of 0.04 dex, the median value for the observed sample. To match the observed sample, we randomly select $20\%$ of sources for which we set [Fe/H] = 0 with an elevated uncertainty of 0.25 dex.
When calculating these absolute magnitudes and their uncertainties, Shahaf et al. (2024) incorrectly used parallaxes and parallax uncertainties from the gaia_source catalog, as opposed to those from the NSS catalog, which accounts for binary motion. While the parallax distribution is minimally affected by this, the reported uncertainties are, on average, a factor of 2.5 higher in the gaia_source catalog. To remain consistent with their method, we therefore increase our mock parallax uncertainty measurements by the same factor in this calculation.
3.1.1 Color excess distribution
Figure 5: Distribution of color excess for systems in the non-class I sample (Left) and those that end up in the final NCE sample (Right). We see that on average, triples have larger color excess and are therefore effectively excluded from the final sample. However, we also find that over half of true WD+MS binaries are removed in the process. Moreover, the distribution for the observed non-class I sample is significantly broader than that of our model, possibly due to inaccuracy in the measured metallicities (Section 3.1.1, Appendix A.3).
Figure 4 shows the reported color excess of the observed non-class I sample as a function of $M_{1}$, [Fe/H], $P_{\rm orb}$, and eccentricity. Recall that Shahaf et al. (2024) set [Fe/H] = 0.0 for $\sim 20\%$ of sources with unreliable metallicites, causing the vertical streak at this value. Several other features stand out. Firstly, we observe a significant anti-correlation between color excess and [Fe/H]: essentially all sources with [Fe/H] $<-0.5$ have significant color excess. The source of this trend is uncertain, with one possibility being that the assumed relationship between color and metallicity is too strong in the PARSEC models (i.e. metal-poor stars are predicted to be bluer than they really are). Metal-poor stars are also older and thus on average are more evolved than they would be at the assumed age of 2 Gyr. This means that the observed sample is likely to be biased against metal-poor binaries. Second, we see there is a correlation between color excess and eccentricity. Hierarchical triples are expected to have higher eccentricities on average than WD+MS binaries, so this trend likely reflects an increasing fraction of triples at high color excess. A further discussion of this trend can be found in Appendix A.1. We also see that there is a cloud of sources with primary masses between $\sim 0.6-1.2\,M_{\odot}$ with systematically negative color excess, suggesting that the PARSEC models predict stars to be redder than they are in reality.
Since we use PARSEC isochrones to calculate both the true and expected colors of the simulated binaries, the color excess distribution of our simulated WD+MS binaries peaks at zero, as expected. To properly reproduce the observed relation with primary mass would require identifying and correcting the underlying cause of the inaccurate colors predicted by the PARSEC models, which we do not attempt to do in this work. Instead, we apply an empirical correction to the color excess of our simulated systems. This is detailed in Appendix A.2. We note that the distributions of key parameters in the observed and modeled populations are comparable prior to the cut on color excess (i.e. the non-class I sample), so details of this correction do not play a critical role in our conclusions. However, the correction does affect the total number of systems that enter the final NCE sample and therefore adds uncertainty to the inferred occurrence rate of wide post-MT systems.
The distributions of color excess for systems that enter the modeled (Section 5) and observed non-class I and final NCE samples are plotted in Figure 5. As expected, we see that on average, triples have significant non-zero color excess which results in the vast majority of them being excluded from the final sample. At the same time, around half of the simulated WD+MS binaries also have non-zero color excess and are excluded, which is primarily due to the fixed age assumption made in the calculation of color excess (Section 3.1) and suggests a less stringent cut may have been possible. In addition, the observed distribution is broader. One possible cause of this is that there is a discrepancy between the true and measured metallicities. This may also be responsible for the difference in the metallicity distributions of the observed and model samples (Figure 7). While there is no obvious cause of this, we note that the Zhang et al. (2023) metallicities were calculated assuming gaia_source parallaxes (as opposed to those from the NSS catalog) and did not account for the presence of light from luminous companions in the XP spectra. Since the uncertainties on these parallaxes are underestimated (El-Badry, 2025), it is likely that they metallicities are less accurate than their uncertainties suggest. The effect of introducing such a discrepancy is explored in Appendix A.3 but it does not significantly affect the distributions of the other parameters in the final NCE sample. However, a broader color excess distribution reduces the total number that enter the sample, adding an uncertainty of $\sim 5\%$ on the inferred fraction of post-AGB MT systems that end up in wide orbits, on top of the that coming from the choice of an empirical correction described above.
3.1.2 Deficit of triples in the simulated population
Figure 6: AMRF against primary mass for the observed (left) and mock (fiducial model; right) non-class I samples. The points are colored by their color excess. The shaded region is the expected location for systems hosting dark secondaries with masses between $0.45-0.75\,M_{\odot}$ (i.e. WD+MS binaries). The dashed and dotted lines are upper limits to systems consistent with hosting a single MS secondary and an inner binary of two MS stars respectively. In both samples, there is a group of sources with near-zero color excess and located in the shaded region corresponding to WD+MS binaries. The others with large positive color excess values are most likely triples, which are rarer in our simulated population than in the observed sample.
In Figure 6, we plot the AMRF against primary mass for sources in the observed and mock (Section 5) non-class I samples, colored by the color excess. The purple strip corresponds to the expected location of systems with dark $0.45-0.75\,M_{\odot}$ companions. The dash and dotted lines are the boundaries below which systems are consistent with hosting a single MS star and an inner binary respectively (i.e. non-class I systems lie above the dotted line; Section 3). These plots can be compared to Figure 5 in Shahaf et al. (2024).
In both samples, there is a high density of sources occupying the strip with color excess close to zero, corresponding to the true WD+MS binaries. The distribution is slightly narrower and shifted to the left for the mock sample, which is reflected in the primary and WD mass distributions as discussed further in Section 6.
Here, we focus on the sources with positive color excess values located outside of the strip and in between the two boundaries, which are presumably the triples. For the simulated sample, the vast majority of triples do indeed fall in this group (Figure 5). However, we see that there are significantly more such likely triples in the observed sample, likely reflecting an imperfect model of compact hierarchical triples in our simulated sample. We could increase the total number of these sources in our simulated population if we alter the mass ratio distribution from which we draw the outer tertiary’s mass so that there are fewer systems for which both stars in the inner binary have evolved into WDs (Section 2.5). Still, this would be unable to reproduce the pileup of observed sources at $M_{1}\sim 0.95\,M_{\odot}$, also seen in Figure 6 of Shahaf et al. (2024). This feature may suggest an additional structure in triple fraction or triple mass ratio distribution for compact triples that is not accounted for in our model. Alternatively, it could be the result of overestimated primary masses for the observed sources in the binary_masses catalog. Since almost all triples are effectively removed by the cut on color excess, these discrepancies have a minimal effect on our final sample but future work to study the nature of these sources may be informative.
4 Results of E24 model
We first consider the original mock NSS catalog from E24, with the addition of hierarchical triples (Section 2.5). This will be referred to as the E24 model. In Figure 7, we compare distributions of several parameters for systems in the observed (filled, light blue) and the simulated E24 (red dashed border) NCE samples. We also plot the results of our new “fiducial” model (purple solid border) where several improvements have been made on the E24 model, as described in Section 5. $M_{\rm 2,dark}$ is the secondary mass implied by the astrometric solution assuming a completely dark companion. In these plots, the total number of systems has been adjusted to be equal for easier comparison of the shapes of distributions. Here, the total number of systems in the mock sample is within $8\%$ of the true sample, suggesting that the assumed $10\%$ fraction of post-AGB MT systems in wide orbits is reasonable, given the other assumptions of the model (see discussion towards the end of Section 6.1).
We see that there is a deficit of systems with $P_{\rm orb}<300\,$d and $e>0.2$ in the E24 model. The color excess cut removes the vast majority of hierarchical triples that enter the non-class I sample so that they make up less than 0.5% of the NCE sample. Therefore, we can conclude that the deficit of systems with $e>0.2$ cannot be attributed to contaminating triples. The distribution of color excess is discussed further for our new fiducial model in Section 6.1 (Figure 5). The lack of shorter period systems can be addressed by modifying the conditions that lead to the formation of wide post-MT systems, as will be described in Section 5.1 and illustrated in Figure 8.
The model significantly overpredicts the fraction of WDs with masses $\gtrsim 0.8\,M_{\odot}$, which we address for our fiducial model in the following sections. This deficit of massive WDs in the AMRF sample compared to the field WD population has been explored recently by Hallakoun et al. (2024) (Section 6.1). Massive WDs are over-represented in the AMRF sample as they produce larger orbits and can therefore be identified as WDs even when they orbit more massive/luminous stars, so the dearth of massive WDs is highly significant.
The observed primary mass distribution is more narrow and peaks at a slightly higher mass compared to the E24 model. This suggests that there may be a relationship between the mass of the accretor in the progenitor system – the luminous primary today – and the probability to end up in wide orbits post-interaction that are detectable via astrometry (Section 5.1).
We also see that there is a slight discrepancy between the observed and simulated [Fe/H] distributions. This may be attributed to uncertainties in the underlying stellar population simulated by Galaxia or those in the measured metallicites reported by Zhang et al. (2023) (Section 6.1).
Figure 7: Distributions of stellar and orbital parameters in the real NCE sample from Shahaf et al. (2024), and mock samples obtained from processing the original mock NSS catalog from E24 (with the inclusion of hierarchical triples; Section 2.5), and our updated fiducial model (Section 5) Overall, the fiducial model improves on the major discrepancies present between the distributions of the E24 model and observed sample.
5 Modifications to the E24 model
The key components of our updated model are summarized in Figure 8, which we describe in the following.
5.1 Orbital evolution under MT
Figure 8: A schematic summarizing the assumed orbital evolution of the zero-age binary population in our fiducial model which lead to the formation short-period and wide post-MT binaries. Detailed descriptions of the components can be found in Section 5.
Figure 9: Maximum radius that a star reaches on the RGB and AGB as a function of its initial mass. In our model, systems with initial pericenter separations below the RGB curve are assumed to become short-period PCEBs, while a fraction of those in between the two curves end up as wide PCEBs.
As mentioned in Section 2.3, the outcome of binary mass transfer is expected to depend on the structure of the donor’s envelope. In particular, compared to red giants, AGB stars can have very loosely bound envelopes that may be ejected with very little energy (e.g. Belloni et al., 2024; Yamaguchi et al., 2024a). We therefore identify and separate the orbital evolution of binaries which interact when the donor is on the RGB versus the AGB by comparing their initial separations to the maximum stellar radii at each of these phases.
We use the standard set of MIST evolutionary tracks (with rotation; Dotter 2016; Choi et al. 2016) to predict the maximum radius reached on the RGB and AGB, $R_{\rm RGB,max}$ and $R_{\rm AGB,max}$ respectively, for a given initial stellar mass (i.e. mass of the WD progenitor). This is plotted in Figure 9. We fit two cubic spline functions for the RGB track above and below $2\,M_{\odot}$. For the AGB track, there is large scatter due to thermal pulses, particularly above $\sim 5\,M_{\odot}$. To avoid over-fitting the data, we excluded local dips in the maximum radii at these masses while fitting a single cubic spline. We note that progenitors with masses above $\sim 5\,M_{\odot}$ make up a small fraction of the total population ($\sim 3\%$) and thus the effect of this choice on the statistics of the final population is small.
We modify the conditions that lead to the formation of short ($P_{\rm orb}=1\,d$) and longer period post-interaction binaries from those described in Section 2. Firstly, instead of using a simple cut on the semi-major axis, we consider the Roche lobe radius at periastron, $r_{\rm L,peri}=f_{q}a_{i}(1-e)$ (where $f_{q}$ is the Eggleton factor; Eggleton 1983). This is reasonable as interaction most likely begins when the two components are closest to each other. We assume that systems which first interact when the WD progenitor is on the first giant branch, where $r_{\rm L,peri}<R_{\rm RGB,max}$, end up with short (111\,1d) periods. Meanwhile, a fraction of systems that first interact on the AGB, where $R_{\rm RGB,max}<r_{\rm L,peri}<R_{\rm AGB,max}$, may form wide post-interaction binaries. In the E24 model, 10% of such systems were randomly chosen to end up in orbits with final separations that are a constant fraction, $\beta=0.5$, of their initial separations. In our updated fiducial model, we instead assume that all systems that interact with an AGB donor evolve with a $\beta$ that depends on the initial mass ratio, $q_{\rm init}=M_{\rm 2,init}/M_{\rm 1,init}=M_{\rm 1,final}/M_{\rm WD,prog}$ such that it increases with higher $q_{\rm init}$. This means that on average, systems with more massive accretors/less massive donors experience less orbital shrinkage and form wider post-MT systems, which are more likely to receive astrometric orbits. This choice is motivated by the discrepancy in the component mass distributions of the final WD+MS binaries in the observed sample and the E24 model (Figure 7). We use a sigmoid function, plotted in Figure 10, which provides a steep dependence on $q_{\rm init}$ required to match the observed deficit of low mass primaries and high-mass WDs. Overall, this results in $13\%$ of post-AGB interaction systems forming “wide” orbits with $P_{\rm orb}=100-1000\,$d. Although a larger fraction of these systems have $q_{\rm init}>0.38$ above which $\beta$ is large (approaching 0.63), the majority end up in periods $>1000\,$d due to the decrease in mass resulting from the formation of the WD. While we ultimately chose a steeply rising sigmoid function that is step-like, in Appendix B, we show the effects of varying the form of this function on the final results.
While it is possible to achieve a similar effect by preserving the E24 setup (leaving $\beta$ constant, instead allowing the fraction that form wide orbits to vary with $q_{\rm init}$), our new choice is more physically motivated. For example, the $\alpha$-formalism prescription for the common envelope phase, which is an argument of energy conservation, predicts a final separation that depends on the component masses and thus implicitly on their ratio. In Section 5.3, we describe an alternative model for $\beta$ built using on the $\alpha$-formalism.
Furthermore, the mass ratio at which the step occurs ($q_{\rm init}=0.38$ in our fiducial model) can be understood as the critical mass ratio (e.g. Ge et al., 2020; Temmink et al., 2023) above which mass transfer remains stable and orbits stay wide, and below which CEE ensues and orbits shrink significantly. As we will discuss further, it is debatable whether stable mass transfer or CEE is a more accurate description of the binary interaction that leads to the formation of wide post-MT binaries. The key input for our model is simply that binaries with extreme mass ratios shrink to small orbits, whereas those with closer-to-equal mass ratios remain relatively wide.
Figure 10: The ratio $\beta$ of the final to initial separation as a function of the initial mass ratio $q_{\rm init}=M_{\rm 2,init}/M_{\rm 1,init}$. The curve is a Sigmoid function whose functional form is given at the bottom right corner of the figure.
We also alter the eccentricity evolution so that the final eccentricity, $e_{f}$, of the wide post-MT systems are related to their initial values, $e_{i}$, by $e_{f}=e_{i}^{4}+0.05$. A gaussian noise with a standard deviation of 0.05 is also added, and an upper bound of $e_{f}=1$ was imposed. This is an ad-hoc prescription which allows us to better reproduce the true distribution. However, simulations of CEE from Glanz & Perets (2021) found a correlation between initial and final eccentricities at the end of the dynamical phase which motivates a choice of this kind relating the two.
5.2 WD IFMR
Figure 11: Core mass as a function of the stellar radius on the AGB for various initial stellar masses. The thicker line weight highlights regions where the radius is greater than at any previous point in the evolution, which are used in the interpolation.
We also adopt a different WD IFMR which takes into account the interruption of the progenitor’s evolution due to the onset of binary interaction. On average, this should lead to lower WD masses compared to those predicted from single star evolution and ameliorate the discrepancy between the observed and simulated WD mass distributions (Section 4). We estimate WD mass as the he_core_mass from the MIST evolutionary tracks at a given giant radius and initial mass. For each system which interacts on the RGB or AGB, the appropriate track corresponding to the closest initial mass is chosen and the core mass at $r_{\rm L,peri}$ is set to be the final WD mass. Figure 11 plots the core mass vs. radius on the AGB for a range of initial masses. Note that during thermal pulses, the growth becomes non-monotonic. However, this is not problematic for our model, as we only consider regions where the radius is larger than at any earlier point in the evolution (otherwise interaction would have occurred earlier). Moreover, we see that the core mass can exceed $1.4\,M_{\odot}$ at the final stages of the AGB phase for progenitor masses $\sim 6-8\,M_{\odot}$. In this case, we return to using the Weidemann (2000) IFMR. This is reasonable as the core is defined in terms of the helium mass fraction and not all of this material ends up burning into carbon and oxygen. A very small fraction of systems are in this regime so the end results are insensitive to this choice.
For stars that interact on the AGB, our adopted IFMR leads to WD masses that are $\sim 10-20\%$ lower than those predicted by the Weidemann (2000) relation for progenitor masses between $\sim 1.5$ and $4.5\,M_{\odot}$. This corresponds to WD masses between $\sim 0.55-0.85\,M_{\odot}$. Outside this range, our IFMR can lead to higher WD masses. For comparison, Shahaf (2025) used a different method on a spectroscopic sample of AU-scale WD + red giant binaries in open clusters and found that these WDs were $\sim 20\%$ less massive than isolated WDs, which they attributed to binary interaction (also see Ironi et al. 2025).
Combined with the changes in the assumed orbital evolution due to MT described above (Section 5.1), we will refer to this modified version of the simulated population as our “fiducial” model. They key components of this model are illustrated in Figure 8. This new simulated population leads to a modified mock NSS catalog which we once again process through the Shahaf et al. (2024) pipeline.
5.3 An $\alpha$-formalism model
The $\alpha$-formalism (Livio & Soker, 1988; de Kool, 1990) is often used to describe the orbital evolution resulting from CEE (e.g. Zorotovic et al., 2010; Davis et al., 2012; Scherbak & Fuller, 2023). Here, we test whether the $\alpha$-formalism can predict the properties of WD+MS binaries formed via interaction on the AGB. At its core, the formalism is a statement of energy conservation:
$E_{\rm bind}=\alpha\left(-\frac{GM_{\rm WD}M_{\star}}{2a_{f}}+\frac{GM_{i}M_{\star}}{2a_{i}}\right)$ (5)
where $E_{\rm bind}$ is the binding energy of the donor’s envelope, and the right-hand side is a fraction $\alpha$ of the orbital energy loss. $M_{i}$ is the initial mass of the WD progenitor (i.e. the donor). In the classic formulation, $E_{\rm bind}=-(GM_{i}M_{i,\rm env})/(\lambda R_{i})$ where $M_{i,\rm env}$ and $R_{i}$ are the envelope mass and radius of the donor, and $\lambda$ is a free parameter which encodes information about the envelope structure, typically set to $\sim 0.5$. Taking $R_{i}$ to be the Roche lobe radius at periastron and $M_{i,\rm env}=M_{i}-M_{\rm WD}$, we can solve for $a_{f}$ with one free parameter, namely the product $\alpha\lambda$.
A value of $\alpha\lambda=1$ is typical under the standard assumptions because $\alpha$ is defined to be less than or equal to 1 and considering only the gravitational binding energy, we expect $\lambda$ to be on the order unity. A higher value may be understood as the donor having a particularly loosely bound envelope, which may be the case if it is undergoing a thermal pulse on the AGB, or if additional sources of energy (e.g. recombination) aid in the unbinding (e.g. Ivanova et al., 2013, 2015; Belloni et al., 2024; Yamaguchi et al., 2024a, a).
We use the same WD IFMR as in Section 5.2 for this model, which we will refer to as the “$\alpha\lambda$ model” hereafter. Although the $\alpha$-formalism is not expected to provide an accurate description of the mass transfer process, we include this model as it may be useful to have an effective value of $\alpha\lambda$ to predict properties of PCEBs without having to use detailed evolutionary models, which themselves struggle to predict the outcome of CEE. The results of using this model can be found in Section 6.3.
6 Results of modified models
6.1 Fiducial model
The distributions of stellar and orbital parameters of the mock NCE sample from our fiducial model are plotted with a solid purple border on Figure 7.
The primary mass distribution of the systems in our fiducial model is more peaked compared to that of the E24 model, in better agreement with the observed distribution. This is the result of introducing a mass ratio dependence to the ratio between initial and final separation (Section 5.1). In this scheme, more massive MS companions are favored to form wide post-MT systems that are detectable by astrometry. Our model peaks at a slightly lower primary mass than the observed distribution. While this could be improved upon by increasing the mass ratio at which the orbital shrinkage decreases (i.e. where the transition occurs in Figure 10), this would lead to a decreased total number of WD+MS binaries in the final sample (Appendix B).
Compared to the E24 model, our model leads to a WD distribution with a clear deficit at high masses, in better agreement with observations. There are two contributing factors. First, our fiducial model takes into account the early termination of the core growth due to binary interaction. While this has a non-negligible effect, just accounting for binary truncation of the core’s mass growth is insufficient to match the observed WD mass distribution. More importantly, the post-MT orbital separation being a function of mass ratio not only affects the luminous star mass distribution, but also significantly reduces the fraction of systems with massive donors (with low $q_{\rm init}$) that form wide orbits, and therefore the fraction of massive WDs in the final sample.
Next, the period distribution of our model also better matches the true distribution compared to the E24 model. Since this is largely determined by the assumed orbital evolution for which we adopt more realistic boundaries of MT from an RGB vs. AGB donor compared to the E24 model (as described in Section 5.1), it suggests that the assumptions made are likely reasonable. It is particularly interesting that only a fraction of MT products from an AGB donor – those with $q_{\rm init}\gtrsim 0.4$ – end up in these relatively wide AU-scale orbits. Since the stellar radii peak during thermal pulses, we might expect interaction to occur most commonly during this phase when the envelope is most loosely bound. But if this is the relevant physical mechanism, our model requires the envelope to be efficiently ejected in only about $5-15\%$ of cases (given the uncertainties arising from the calculation of color excess; Section 3.1.1). If AGB mass transfer always resulted in a wide orbit, the model would significantly overpredict the observed number of WD+MS binaries.
Although our modeling is agnostic to whether AGB mass transfer is more akin to stable mass transfer or CEE, the mass ratio-dependent $\beta$ assumed in our model could also naturally be interpreted as a criterion for mass transfer stability: that is, orbits remain wide if the accretor-to-donor mass ratio exceeds $\sim 0.4$, and otherwise shrink significantly.
By construction, the eccentricity distribution now closely matches that of the real sample. We emphasize our simulations self-consistently account for the eccentricity bias in the astrometric orbits (e.g. Wu et al., 2024), and still we require an underlying eccentricity distribution that peaks around $0.1$ to reproduce the observed eccentricities of WD+MS binaries in astrometric orbits. While some previous works have suggested that these non-zero eccentricities are a result of triple dynamics (e.g. Belloni et al., 2024), we consider it highly unlikely that all eccentric binaries – a large majority of the sample – have unseen tertiaries.
6.2 Intrinsic properties of WD+MS binaries
In the previous section, we made changes to the simulated binary population which was then processed by the NSS (Halbwachs et al., 2023) and AMRF (Shahaf et al., 2024) analysis pipelines to produce a mock NCE sample whose parameter distributions approximately match the Gaia WD+MS binary sample. We are now in a position to study the properties of the underlying WD+MS binary population.
6.2.1 Completeness fractions
Figure 12: (A): Completeness fraction of the fiducial mock NCE sample binned by orbital period and grouped by distance ranges, for our fiducial model. As expected, the completeness falls at further distances. The dip in completeness at $\sim 1\,$yr is due to the confusion between parallactic and orbital motion. (B): Completeness fraction binned by orbital period at different stages in the processing pipeline for sources within 500500500\,500pc. (C): Same as panel B, but binned by primary mass. We see that while completeness for the NSS catalog generally rises for more massive primaries as they are brighter on average, they are disfavored by the cut on AMRF. (D): Same as panel B, but binned by WD mass. While there is a bias towards more massive WDs for a given primary mass for all processing steps, the positive correlation between the two complicates this.
We first calculate the completeness fraction as a function of a few different parameters. We define completeness as the number of systems that make it into the final NCE sample in each parameter bin divided by the total number of all WD+MS binaries in the simulated population in the same bin. For the orbital period, we take bins of width 505050\,50d from $100$ to 100010001000\,1000d, and consider sources in several distance ranges. The resulting completeness fractions are plotted for our fiducial model in panel A of Figure 12. In panels B, C, and D we plot the completeness fractions at different steps in the processing pipeline (for sources at distances $<500\,$pc) binned by period, primary mass, and WD mass, respectively.
The dip at $\sim 1\,$yr is a selection effect that results from confusion between parallactic and orbital motion. This prevents systems with orbital periods close to 1 yr from entering the NSS catalog, and any of its products (including the non-class I and NCE samples). As expected, the completeness drops at all periods as we move out to larger distances. The total number of systems in the simulated population is greatest below 505050\,50d but due to their small orbits, almost none of them receive orbital solutions in the NSS catalog, making their completeness close to zero.
Moreover, we see that completeness of the NSS catalog increases with primary mass as more massive primaries are, on average, more luminous. Meanwhile, there is a sharp drop off beyond $\sim 0.8\,M_{\odot}$ for the non-class I sample, since the AMRF cuts disfavor massive primaries. In addition, more massive primaries evolve faster and are thus redder on average, making them more likely to be removed by the cut on color excess.
Lastly, we find that the completeness is generally higher for more massive WDs in the NSS catalog as they will produce wider orbits at fixed period, which are easier to detect. The effect of the AMRF cut on the completeness is more complicated, because while more massive WDs are favored for a given primary mass, they are more likely to be found around more massive and luminous primaries. These trends are described further in Section 6.2.2.
6.2.2 Parameter distributions
Figure 13: Distributions of orbital period and component masses of WD+MS binaries, with periods between $100-1000\,$d and at distances within 500 pc, at different stages in our analysis: the full simulated WD+MS binary population, systems that enter the mock NSS catalog, those in the non-class-I sample, and lastly, those in the NCE sample. In magenta, we have the intrinsic period distribution inferred by taking the distribution of the observed NCE sample and correcting it with the completeness implied from our models (Figure 12). We see that the intrinsic period distribution is almost flat over $100-1000\,$d, rising gradually toward shorter periods.
Figure 14: Total number of simulated WD + MS binaries with $P_{\rm orb}=100-1000\,$d within 500 pc at each step in the processing pipeline.
In Figure 13, we plot distributions of the orbital period and component masses of the intrinsic WD+MS binary population with $P_{\rm orb}=100-1000\,$d, the majority of which are post-interaction systems. We also add plots of the resulting mock NSS catalog and the mock non-class I and NCE samples. The shaded region around the intrinsic period distribution represents the uncertainties estimated as the ratio of the square root of the number of systems in the NCE sample (i.e. Poisson error) over the completeness fraction in each period bin (Section 6.2.1). In Figure 14, we provide a flowchart showing the total numbers of WD + MS binaries that make it through each step of the pipeline.
The observed orbital period distribution largely reflects the completeness fraction. Once again, there is a dip at $\sim 1\,$year and a drop towards shorter periods; these features are a result of the lower sensitivity of astrometry to these orbits. The intrinsic distribution has a shallow power law slope, $\mathrm{d}N/\mathrm{d}P_{\rm orb}\propto P_{\rm orb}^{-0.54\pm 0.01}$ (the error in the slope is calculated from the residuals to the fit), increasing towards shorter periods. Relative to the intrinsic population, there are fewer low-mass luminous stars in the NSS catalog. This reflects the fact that low-mass stars are faint and only yield high-quality orbits at close distances. As discussed in Section 6.1, cuts based on the AMRF and color excess also preferentially remove higher-mass primaries. This is because more massive stars require higher mass companions to eliminate the possibility of a luminous secondary. They also evolve more quickly, making them appear redder than expected assuming a fixed age. Few WDs at the lowest masses $\lesssim 0.5\,M_{\odot}$ enter the final sample. One reason for this is that for a given primary mass, lower mass WDs will produce smaller astrometric wobbles and are thus less likely to receive orbital solutions. Moreover, for a system with a given primary mass, those with low mass WD companions are less likely to be unambiguously identified as having non-luminous companions (i.e. most will not make it into the non-class I sample). However, as there is a positive correlation between the masses of components in the zero-age binary population and massive primaries are disfavoured by the AMRF cut, this last effect evidently does not dominate at high WD masses.
In the leftmost panel, in magenta, we plot the period distribution of the observed NCE sample divided by the completeness fraction inferred using our mock NCE sample. As expected, this closely matches the “intrinsic” curve, though it is slightly flatter with $\mathrm{d}N/\mathrm{d}P_{\rm orb}\propto P_{\rm orb}^{-0.25\pm 0.10}$. There is some deviation at the shortest ($\lesssim 100\,$d) and longest ($\gtrsim 900\,$d) periods which simply reflects the inconsistencies between our model and observed NCE samples as seen on bottom left panel of Figure 7.
In total, in our simulated binary population, there are $242$ WDs around solar-type stars (which are taken to have masses between $0.8-1.2\,M_{\odot}$) with $P_{\rm orb}=100-1000\,$d, located within 100 pc. There are $\sim 330,000$ Gaia sources within this volume (which is almost complete; Gaia Collaboration et al. 2021). From this, selecting those with extinction-corrected G band absolute magnitudes between $3.0-6.2$ and BP-RP between $0.6-1.1$ (based on MIST isochrones for MS stars of masses $0.8$ and $1.2\,M_{\odot}$), we find that $\sim 8.2\%$ are solar-type stars. Given this, we calculate that $\sim 0.9\%$ of solar-type stars have WD companions at these orbital periods. This is roughly consistent with the rate of these systems inferred from WD self-lensing binaries discovered with Kepler (Masuda et al., 2019; Yamaguchi et al., 2024b).
6.3 $\alpha\lambda$ model
Here, we discuss predictions of the $\alpha\lambda$ model (Section 5.3), where the $\alpha$-formalism prescription is used to describe the orbital shrinkage resulting from a common envelope phase with an AGB donor.
Firstly, setting $\alpha\lambda=1$ corresponds to $\beta\lesssim 0.1$ (where $\beta$ is the ratio of the final over initial separation; Section 5.1). This results in very few systems entering the non-class I and thus final NCE sample. This issue has already been explored by several recent works (e.g. Yamaguchi et al., 2024a, b; Belloni et al., 2024) – the AU-scale WD+MS binaries discovered by Gaia cannot be explained as a result of CEE with $\alpha\lambda\sim 1$ (Zorotovic et al., 2010; Davis et al., 2012; Toonen & Nelemans, 2013; Scherbak & Fuller, 2023). Instead, they require for the donor’s envelope to be almost unbound. In Figure 15, we plot the results of setting $\alpha\lambda=10^{4}$, which corresponds to a range of $\beta\sim 0.15-0.65$. In this case, the model predicts about 10 times more WD+MS binaries in the final sample than are observed. To remedy this tension, we assume that only 10% of systems that interact on the AGB have $\alpha\lambda=10^{4}$, while the remainder have $\alpha\lambda=1$ and form short-period PCEBs. We see that the model predicts proportionally too many systems with $P_{\rm orb}\lesssim 1\,$yr and too few with $P_{\rm orb}\gtrsim 750\,$d. This is because the intrinsic period distribution of WD+MS binaries produced by this model rises more rapidly towards shorter periods compared to our fiducial model (Figure 13). Further increasing $\alpha\lambda$ does not alleviate this problem as $\beta$ asymptotes to the ratio of the WD mass over its progenitor mass as $\alpha\lambda\rightarrow\infty$. Moreover, this model does not reproduce the highly peaked WD mass distribution. This is because at sufficiently large values of $\alpha\lambda$, $\beta$ depends weakly on the initial mass ratio of the system. Given these limitations, we conclude that the $\alpha$ formalism is not well suited to describe the formation of this wide WD+MS binaries. All of the same issues exist for a more modest value of $\alpha\lambda=100$, but with there being an even more disproportionate abundance of systems with $P_{\rm orb}\lesssim 1\,$yr.
Figure 15: Analogous to three panels from Figure 7 with the same x axes, but showing the results of using the $\alpha$-formalism to evolve orbits as described in Section 5.3. Here, we set $\alpha\lambda=10^{4}$, which lead to wide enough orbits that that can be detected via astrometry. Overall, there are larger discrepancies between the distributions of the observed sample and this model compared to our fiducial model.
7 Conclusion
Using orbital solutions from the Gaia NSS catalog, Shahaf et al. (2024) assembled a sample of over 3000 high-probability WD+MS binary candidates in AU-scale orbits. The formation histories of these systems are not well understood, motivating a population-level study to better understand their occurrence rates and population demographics. However, this requires an understanding of the selection function that results from both the construction of the Gaia NSS catalog and the cuts employed by Shahaf et al. (2024) to select WD+MS binary candidates.
In this work, we use a forward modeling approach to characterize the underlying population of WD+MS binaries, extending the work of E24. We summarize the key findings below:
  • We make modifications to the E24 model for the evolution of binaries after binary interactions. In particular, for our fiducial model, we use MIST stellar models to impose more realistic bounds for interaction with donors on the RGB or AGB, and to account for the early termination of core growth in the WD initial-final mass relation. We also let the ratio of the final to initial separation depend steeply on the initial mass ratio of the system, such that orbits remain wide above a critical value of $M_{\rm accretor}/M_{\rm donor}\sim 0.38$ and shrink significantly below it. (Section 5)
  • While a significant number of hierarchical triples may enter the non-class I sample, the cut on color excess removes the vast majority of them and therefore, we expect the final WD+MS binary sample selected by Shahaf et al. (2024) to be pure. However, our modeling predicts that about half of WD+MS binaries that receive orbital solutions are excluded by the color excess cut. We also find that this cut likely causes the observed NCE sample to be biased against metal-poor systems. A less stringent selection, perhaps one with a dependence on metallicity, may be implemented so that fewer true WD+MS binaries are removed.
  • The simulated sample of astrometric WD+MS binaries obtained from our fiducial model broadly reproduces the observed period and component mass distributions (Figure 7). $28\%$ of systems in our mock NSS catalog are WD + MS binaries, out of which $\sim 7\%$ enter the final NCE sample.
  • Our model predicts that $5-15\%$ of stars that interact with an AGB donor stay in wide orbits. A significantly higher fraction is ruled out as this would lead to too many wide (i.e. AU-scale) post-interaction systems (Section 5.1).
  • In our fiducial model, the binaries that do not remain wide are those which have small accretor-to-donor mass ratios. A steep dependence of the amount of orbital shrinkage on the initial mass ratio is required to produce the strongly peaked WD mass distribution seen in the observed sample (Section 6). Such a relation can be physically motivated by the concept of the critical mass ratio that defines the boundary between stable and unstable MT.
  • The post-MT eccentricities of AU-scale WD+MS binaries are typically between 0 and 0.2. Completeness does not significantly affect this distribution. (Section 5.1, Figure 7)
  • Using the $\alpha$-formalism prescription to evolve orbits, we require very large $\alpha\lambda\sim 10^{4}$ for $\sim 10\%$ of systems that interact with an AGB donor to form sufficiently wide WD+MS binaries. However, the final sample resulting from this model overpredicts the proportion of short period ($P_{\rm orb}\lesssim 1\,$yr) systems and does not reproduce the deficit of high-mass WDs (Figure 15, Section 6.3). Moreover, it is unclear why such a high $\alpha\lambda$ value would apply to only $\sim 10\%$ of AGB donors. We conclude that the formation of wide WD+MS binaries cannot be easily understood through this formalism.
  • The completeness of the Gaia WD+MS binary sample as a function of orbital period shows a characteristic dip at 1 yr due to the degeneracy between parallactic and orbital motion and decreases towards the outskirts at $\sim 100$ and 100010001000\,1000d. Taking this into account, we find that the intrinsic orbital period distribution of WD+MS binaries between this period range has a shallow slope, where $\mathrm{d}N/\mathrm{d}P_{\rm orb}\propto P_{\rm orb}^{-0.25\pm 0.10}$.
  • The completeness varies non-trivially with the mass of the primary and WD (Figure 12). This is because while more massive primaries are on average more luminous and thus preferentially receive orbital solutions, they are disfavored in ruling out luminous secondaries. Meanwhile, at a given primary mass, there is a bias towards more massive WDs but this is offset by the intrinsic correlation that exists between the component masses.
  • Our model implies that around $1\%$ of all solar-type stars have WD companions in with $P_{\rm orb}\sim 100-1000\,$d, in agreement with constraints from the observed number of WD self-lensing binaries.
8 Acknowledgements
This research was supported by NSF grant AST-2307232.
NY acknowledges support from the Ezoe Memorial Recruit Foundation scholarship.
This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.
Appendix A Color excess
A.1 Relation between triple fraction and metallicity
Figure 16: Eccentricity distribution of the Shahaf et al. (2024) non-class I sample grouped by [Fe/H]. The ranges are chosen such that there is roughly the same total number in each group. For [M/H] $>$ -0.1, there is an excess of low-eccentricity systems, likely due to WD+MS binaries which have undergone some circularization. For lower metallicities, the eccentricities are higher, possibly reflecting a higher fraction of triples.
In Section 3.1, we noted the presence of a correlation between color excess and eccentricity in the observed non-class I sample (Figure 4). Since we generally expect triples to have higher eccentricities than WD+MS binaries, this may be attributed to a higher triple fraction at large color excess. Figure 4 also shows a strong anti-correlation between color excess and metallicity. This could reflect (a) deficiencies in the color excess metric that are more important for old and metal-poor stars, (b) an increased triple fraction at low metallicity, or (c) a bias in the measured metallicities for triples. To discriminate between these possibilities, we investigate the eccentricity distributions of sources in different color excess ranges.
In Figure 16, we plot the eccentricity distribution of sources in the non-class I sample binned in several different metallicity bins. The eccentricity distributions of the more metal-poor samples are broader than those of metal-rich samples, with more eccentric systems making up a larger portion of the population. This suggests that the trend in color excess with metallicity is not simply a result of biased color excess measurements at low metallicity, as in this case one would expect a low-eccentricity population to also be present at low metallicities. A simpler explanation is that the presence of light from unmodeled secondaries and tertiaries may bias the metallicity measurements from Zhang et al. (2023) toward lower metallicities. It is also possible that compact triples are simply more common at low metallicity. We defer further exploration of the metallicity – color excess anti-correlation to future work.
A.2 Empirical color excess correction to models
We apply an empirical correction to the color excess of our modeled systems to reproduce the observed trend between color excess and primary mass. In theory, true WD+MS binaries are expected to have color excess values close to zero, with more scatter toward positive values due to systems in which the MS star is older than 2 Gyr and has evolved significantly. Therefore, the observed offset towards negative color excess is indicative of a mismatch between the observed colors and those predicted by PARSEC isochrones. We fit the observed color excess values for sources with $M_{1}=0.7-1.2\,M_{\odot}$ and color excess ¡ 0.1 (left panel of Figure 17) with a Gaussian. To reproduce the observed scatter, we also add Gaussian noise with a standard deviation of 0.05 to the simulated color excess values. The fit to the observed non-class I sample is shown on panel A of Figure 17. Panels B and C show the systems in our fiducial model before and after the correction, respectively.
Figure 17: 2D binned plots of color excess against the luminous primary mass of the non-class I systems. From left to right, we have the (A) observed Shahaf et al. (2024) sample with the fitted Gaussian curve used for the empirical correction (Section A.2) (B): our fiducial model with no corrections made in the calculation of color excess, (C): same as (B) but with the empirical correction on color excess applied, and (D): same as (C) but with metallicities drawn from the observed Zhang et al. (2023) distribution (Section A.3).
A.3 Discrepancy in the color excess distribution
As seen in Figure 5, the distribution of color excess is significantly more peaked around zero for our model compared to the observed sample. A potential cause is that the measured metallicities either (a) do not reflect the true values or (b) are drawn from a broader distribution than assumed in our Galaxia simulations. Here, we test this possibility by randomly drawing metallicities and their uncertainties from the observed non-class I sample in calculating $(B-I)_{\rm expected}$. Similarly to our fiducial model, we once again select 20% of the systems to have solar metallicity with an error of 0.25 dex. In Figure 18, we show the color excess distributions of the non-class I and NCE samples obtained after this change has been made. The mismatch introduced between the true and “measured” metallicities significantly broadens the color excess distribution compared to that in Figure 5 so that it is in better agreement with the observed sample. The broadening is also seen in panel D of Figure 17, which shows a 2D histogram of color excess and primary mass. We conclude that biases in the inferred metallicities for triples could plausibly explain both the broadness of the observed color excess distribution and the observed trend in color excess with metallicity. This possibility is also appealing because the metallicity distribution of the observed non-class I sample is significantly broader, and centered on more metal-poor values, than the metallicity distribution of all stars in the solar neighborhood (e.g. Hayden et al., 2015). However, we emphasize that this has minimal effect on the distributions of key stellar and orbital parameters in the final WD+MS binary sample plotted in Figure 7, because the triples with potentially biased metallicities are ultimately excluded from the final WD+MS binary sample.
Figure 18: Analogous to Figure 5, but with metallicities used to calculate $(B-I)_{\rm expected}$ for the modeled systems drawn from the Zhang et al. (2023) metallicity distribution of the observed non-class I sample. As expected, the discrepancy introduced between these mock “measured” metallicities and true values used to calculate $(B-I)_{\rm observed}$ leads to broader distributions (c.f. Figure 5) which better match the observed sample.
Appendix B Varying the $\beta$ function
As described in Section 5.1, our fiducial model assumes that systems that interact with an AGB donor evolve such that the final separation is a fraction $\beta$ of the initial separation which is a function of the initial mass ratio (Figure 10). This is chosen to be sigmoid function with the general functional form:
$\beta\left(q_{\rm init}\right)=\frac{A}{1+e^{-k(q_{\rm init}-q_{0})}}$ (B1)
where $A$ determines the asymptotic value to which the function maximizes, $k$ controls the steepness, and $q_{0}$ defines the critical mass ratio about which beta increases and the orbit remains wide. In our fiducial model, we set $A=0.63$, $k=60$, and $q_{0}=0.38$ (Figure 10). Here, we show the effects of varying each of these three parameters on the final distributions of the orbital period and component masses. On the leftmost panel of Figure 19, we plot the three additional $\beta$ functions tested. To the right, we show the parameter distributions of the final NCE samples obtained using these models. Once again, in order to more easily compare the shapes of the distributions, in these plots, the total numbers of sources in all samples have been adjusted to be equal.
We see that increasing the value of $A$ (i.e. reducing the amount of orbital shrinkage) proportionately increases the number of systems at shorter periods. This is because more systems in initially smaller orbits stay wide enough to receive orbital solutions. The total number of systems that enter the final sample is comparable to the fiducial model, and thus to the observed sample. The component mass distributions are also minimally changed from the fiducial model.
Reducing $k$, which makes the rise in $\beta$ shallower, leads to an orbital period distribution more steeply increasing towards shorter periods. As systems across a wider range of initial mass ratios remain in wide orbits and receive orbital solutions, the distributions of the component masses are slightly broader. This also leads to too many systems entering the final sample, with roughly three times more WD+MS binaries in the final sample than are in the observed sample.
Lastly, pushing $q_{0}$ to larger values primarily affects the component masses. As expected, the peak of the primary mass distribution occurs at a slightly larger value compared to the fiducial model and the deficit of high mass WDs is even more pronounced. As discussed in Section 6, this significantly reduces the total number of systems in the final sample, in this case to about half of that in our fiducial model, because systems must have higher initial mass ratios to end up in wide orbits.
Figure 19: The leftmost panel plots the $\beta$ functions tested. Each of the three additional models have one free parameter modified from our fiducial model (Section 5; Figure 10). The three panels to the right are analogous to those in Figure 15, but for the resulting final NCE samples obtained from using the modified $\beta$ functions. Note that in these plots, the total numbers have been adjusted to be the same across the models. Compared to our fiducial model, the modified versions perform worse in reproducing the observed period distribution and/or lead to too many or few systems entering the final sample.